OpenStudy (nuccioreggie):

An equation is shown below: 5x + 3(x – 5) = 6x + 2x – 15 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation. (4 points)

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

5x+3(x-5)=6x+2x-15 5x+3x-15=6x+2x-15 (distribute) 8x-15 = 8x -15 (combine like terms) 0=0 There are no solutions. I used the Distributive Property when I multiplied 3(x-5)